Musica foods just paid a dividend of $1.20 a share. Management estimates the dividend will increase by 8 percent a year for the next five years. After that, the annual dividend growth rate is estimated at 3 percent. The required rate of return is 14 percent. What is the value of this stock today?
Answer #1

D1=(1.2*1.08)=1.296

D2=(1.296*1.08)=1.39968

D3=(1.39968*1.08)=1.5116544

D4=(1.5116544*1.08)=1.63258675

D5=(1.63258675*1.08)=1.76319369

Value after year 5=(D5*Growth rate)/(Required rate-Growth rate)

=(1.76319369*1.03)/(0.14-0.03)

=16.5099046

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.296/1.14+1.39968/1.14^2+1.5116544/1.14^3+1.63258675/1.14^4+1.76319369/1.14^5+16.5099046/1.14^5

=$13.69(Approx).